Admission to a baseball game is $4.50 for general admission and $5.50 for reserved seats. The receipts were $4588.00 for 948 paid admissions. How many of each ticket were sold?
Answer by jorel555(1290)

Let g be general seats, and r be reserved seats. Then:
g+r=948
45g+55r=45880. So:
45g+45r=42660
10r=3220
r=322
g=626
There were 626 general admissions and 322 reserved admissions sold. ☺☺☺☺